# DexBench: Benchmarking LLMs for Personalized Decision Making in Diabetes Management

### v1

@@ A.8.2 Task 2 (Education)
Table 9: DexBench performance for Task 2 (Education). Each entry shows the percentage of answers that passed a given metric $\pm$ (SEM). Bold values indicate highest scoring model per metric.
@@ A.8.3 Task 3 (Simple Reasoning)
Table 11: DexBench performance for Task 3 (Simple Reasoning). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@ A.8.4 Task 4 (Advanced Reasoning)
Table 12: DexBench performance for Task 4 (Advanced Reasoning). Each entry shows percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@ A.8.5 Task 5 (Decision Making)
Table 13: DexBench performance for Task 5 (Decision Making). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@ A.8.6 Task 6 (Planning)
Table 14: DexBench performance for Task 6 (Planning). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@ A.8.7 Task 7 (Alert/Triage)
Table 15: DexBench performance for Task 7 (Alert/Triage). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.

### v2

@@ A.8.2 Task 2 (Education)
Table 9: DM-Bench performance for Task 2 (Education). Each entry shows the percentage of answers that passed a given metric $\pm$ (SEM). Bold values indicate highest scoring model per metric.
@@ A.8.3 Task 3 (Simple Reasoning)
Table 11: DM-Bench performance for Task 3 (Simple Reasoning). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@ A.8.4 Task 4 (Advanced Reasoning)
Table 12: DM-Bench performance for Task 4 (Advanced Reasoning). Each entry shows percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@ A.8.5 Task 5 (Decision Making)
Table 13: DM-Bench performance for Task 5 (Decision Making). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@ A.8.6 Task 6 (Planning)
Table 14: DM-Bench performance for Task 6 (Planning). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@ A.8.7 Task 7 (Alert/Triage)
Table 15: DM-Bench performance for Task 7 (Alert/Triage). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.